Maria bicycles 6km an hour faster than Praedup. In the same time it takes Praedup to bicycle 42km, Maria can bicycle 60km. How fast does each bicyclist travel?
Answer by checkley79(3341)

D=RT
60=(X+6)T MARIA'S RIDE
T=60/(X+6)
42=XT PRARDUP'S RIDE
T=42/X
60/(X+6)=42/X
42(X+6)=60X
42X+262=60X
42X-60X=-262
-18X=-262
X=-262/-18
X=14 KPH FOR PRAEUR'S SPEED.
14+6=20 KPH FOR MARIA'S SPEED.
PROOF:
60/20=42/14
3=3